# BlackboxBench: A Comprehensive Benchmark of Black-box Adversarial Attacks

### v1

@@ III-A2 Score-based query attacks
The term $\mathbb{B}_{\boldsymbol{x},\epsilon}$ denotes a $\ell_{p}$ norm ball certered at $\boldsymbol{x}$ with radius $\epsilon$. $\xi$ denotes the distance function which returns 0 if $x^{*}$ locates in the ball $\mathbb{B}_{\boldsymbol{x},\epsilon}$ otherwise returns $\infty$. It ensures the imperceptibility of adversarial perturbations. The variable $\Delta$ differentiates based on the type of attack:
@@
The margin loss, captured by the value of $\Delta$, should always be non-negative. A value of 0 for this loss means the attack successfully found its adversarial example, suggesting that the attack process terminates. The primary objective of the attack is to find successful adversarial perturbations while maintaining them within the $\ell_{p}$ norm ball $\mathbb{B}_{\boldsymbol{x},\varepsilon}$. The graphical illustration of untargeted score-based attacks is shown in Fig. 2(b).
@@ III-A3 Transfer-based attacks
Unlike query-based attacks that demand repeated queries, transfer-based attacks are inspired by the phenomenon that adversarial samples produced to mislead a surrogate model $f^{\prime}$ can mislead target models $f$, even if their architectures are significantly different[66], as long as $f$ and $f^{\prime}$ solve the same tasks. Let $\mathcal{L}$ be the loss function (e.g. cross entropy). To fool the classifier $f$, transfer-based attacks use gradient-based methods to craft $\ell_{p}$ norm bounded adversarial perturbation $\boldsymbol{\delta}_{\boldsymbol{x}}$ against the white-box models $f^{\prime}$,
@@
The graphical illustration of untargeted transfer-based attacks is shown in Fig. 2(c).
III-B Implemented Algorithms
Black-box Adversarial Attack Methods Query-basedScore-based Random Search SimBA[7], Parsimonious attack[8], Square attack[9], PPBA[10], BABIES[17] 111111111 Gradient Estimation NES[11], Bandits[12], ZO-signSGD[13], SignHunter[14], AdvFlow[15], NP-attack[16] 111111111 Combination 111111111 $\mathcal{CG}$-attack[18], MCG[19] 111111111 111111111 111111111 111111111 111111111 Decision-based Random Search Boundary attack[92], Evolutionary attack[93], GeoDA[94], SFA[95], Rays[96] 111111111 Gradient Estimation OPT[97], Sign-OPT[98], Triangle attack[99], HSJA[100], QEBA[4], NonLinear-BA[5], PSBA[6] Transfer-based| Data Perspective PGD[21], TI-FGSM[22], SI-FGSM[23], Admix[24], DI2-FGSM[25] 111111111 111111111 Optimization Perspective MI-FGSM[26], NI-FGSM[23], PI-FGSM[27], VT[28], RAP[29], LinBP[30], SGM[31] Feature Perspective ILA[32], FIA[33], NAA[34] 111111111 111111111 111111111 111111111 111111111 Model Perspective RD[35], GhostNet[36], DRA[37], IAA[38], LGV[35], SWA[39], Bayesian attack[39] Composite Attack MI-DI, MI-DI-TI, MI-DI-TI-SI, VMI, VNI, SI-RAP, LGV-GhostNet 111111111 111111111 111111111
We present an elaborate taxonomy of black-box adversarial attacks in Fig. 3. The first level of our taxonomy distinguishes three types of black-box adversarial attacks: decision-based query attacks, score-based query attacks, and transfer-based attacks. At the second level, we subdivide the existing literature within each type based on their technical routes. At the third level, we list the implemented attacks in our BlackboxBench. Given the division of the first level introduced in Sec. III-A, in this section, we introduce our categorization at the second level and the algorithms incorporated into BlackboxBench for each category at the third level. Regarding the method selection, we include classical techniques as baselines and advanced methods to underscore recent advancements in this field.
III-B1 Categorization of implemented decision-based attacks
In the realm of this research, the focus primarily lies on formulating effective strategies that progressively move the adversarial point toward the decision boundary in each iteration. Regarding the methodologies employed for decision-based attacks, they predominantly bifurcate into two core categories for acquiring the gradient to update the current adversarial examples:
  • Random search methods These methods involve random direction updates and heuristic parameter adjustments. They include diverse strategies to acquire updating directions, such as sampling from normal distributions, using evolutionary algorithms, gradient-free iterative techniques, and discrete optimization. (e.g., Boundary attack[92], Evolutionary attack[93], GeoDA[94], SFA[95] and Rays[96]).
  • Gradient-estimation-based methods These methods rely more on the information from the model feedback to guide the search direction. They transform decision-based attacks into optimization problems and incorporate gradient estimation, step size calculation, and binary searching as a framework. Subsequent methods enhance gradient estimation within this framework, and others introduce geometric patterns for improved adversarial detection. (e.g., OPT[97], Sign-OPT[98], HSJA [100], QEBA[4], NonLinear-BA[5], PSBA[6] and Triangle attack[99]).
III-B2 Categorization of implemented score-based attacks
Score-based attacks typically initiate from a point within the $\ell_{p}$ ball and progressively move the sample closer to the decision boundary. The objective of the attack is to generate adversarial samples that reside as close as possible to the decision boundary, allowing minimal input perturbations to mislead the model’s classification. Attackers utilize the predicted scores of the target model to guide the movement direction. These methods are usually classified into three categories:
  • Random search methods These methods usually use stochastic strategies (e.g., sampling) to iteratively modify adversarial perturbations (e.g., SimBA[7], Parsimonious attack[8], Square attack[9], PPBA[10] and BABIES[17]).
  • Gradient-estimation-based methods These methods utilize model feedback or prior knowledge to approximate gradients for adversarial example creation. (e.g., NES[11], Bandits[12], ZO-signSGD[13], SignHunter[14], AdvFlow[15] and NP-attack[16].
  • Combination-based methods These methods enhance adversarial example generation efficiency by integrating extra information from surrogate models, resulting in fewer queries. (e.g., $\mathcal{CG}$-Attack[18], MCG[19]).
As a result, we implemented 25 query-based black-box attack methods in total, covering two types of query feedback, decision and score, returned by the attacked model. 13 decision-based attacks in BlackboxBench are summarized as two categories. 12 score-based attacks in BlackboxBench are summarized as three categories. The general idea of each method will be presented in Appendix D-B and D-A.
III-B3 Categorization of implemented transfer-based attacks
Amongst the earliest transfer-based attacks was the Iterative Fast Gradient Sign Method (I-FGSM)[20] which optimizes Eq. 5 iteratively, serving as a stronger adversary for white-box models compared to the single-step attack—Fast Gradient Sign Method (FGSM)[107]. However, iterative attacks tend to overfit the surrogate model and thus cannot transfer across models. So can we achieve better trade-off between attack ability and transferability? According to a general formulation:
$\arg\max_{\boldsymbol{x}^{*}}\mathcal{L}\left(f^{\prime\prime}\left(\mathcal{T}\left(\boldsymbol{x}_{t}^{*}\right);\theta^{\prime\prime}\right),y;\textbf{I}\right)\text{, s.t. }\left\|\boldsymbol{x}^{*}-\boldsymbol{x}\right\|_{p}\leq\epsilon,$ (7)
  • Data perspective Methods from this perspective use input transformation $\mathcal{T}\left(\cdot\right)$ on the benign sample to alleviate the overfitting to the surrogate model. (e.g., PGD[21], TI-FGSM[22], SI-FGSM[23], Admix[24], DI2-FGSM[25]).
  • Optimization perspective Methods from this perspective introduce gradient-based optimization algorithms to escape from poor local maxima or modify the gradients to facilitate transferability. (e.g., MI-FGSM[26], NI-FGSM[23], PI-FGSM[27], VT[28], RAP[29], LinBP[30], SGM[31]).
  • Feature perspective Methods from this perspective distort intermediate layer features by designing a new adversarial loss function $\mathcal{L}\left(\cdot;\textbf{I}\right)$ with the aid of auxiliary information I (e.g., ILA[32], FIA[33], NAA[34]).
  • Model perspective Methods from this perspective adjust the basic surrogate model $f^{\prime}$ to $f^{\prime\prime}$ (parameterized by $\theta^{\prime\prime}$) to improve transferability (e.g., RD[35], GhostNet[36], DRA[37], IAA[38], LGV[35], SWA[39], Bayesian attack[39]).
As shown in Fig. 3, apart from I-FGSM, BlackboxBench incorporates 22 additional transfer-based attacks falling into these four categories. Note that the category Composed attack in Fig. 3 indicates compositions of aforementioned individual transfer methods, aimed at harnessing the strengths of each approach for maximum benefit. 7 composite attacks are designed officially in BlackboxBench but more composites could be customized easily by flexibly combining existing functional blocks to constitute the attack pipeline. More detailed descriptions of transfer-based attacks are presented in the Appendix D-C.
III-B4 Categorization of implemented defenses
  • Training-time-based approaches These defenses focus on enhancing the robustness of the classifier itself. It requires significant data augmentation during training, which can be costly and typically isn’t easily adjustable during inference. The most prominent training-time defense is Adversarial Training (AT) [65, 2, 1, 91, 60, 58, 59], which is a defense strategy designed to bolster the robustness of machine learning models against adversarial attacks by incorporating adversarial examples during training. While this method enhances resilience, it demands more computational resources and may compromise accuracy on clean data. Typically, adversarial examples are generated using gradient-based techniques, and the training not only aids in generalizing against diverse attack methods but also offers resistance to transfer attacks. Despite its effectiveness, there’s a trade-off between robustness and clean data accuracy, and various adaptations of adversarial training, like ensemble versions, have emerged to optimize its performance. Essentially, adversarial training proactively modifies model parameters for increased robustness, distinguishing it from static defenses that focus on input or data alterations without changing the model itself. AT is implemented in BlackboxBench and deployed against both query-based attacks and transfer-based attacks.
  • Inference-time-based approaches These defenses are generally more lightweight since they are applied only during the inference of the classification. Compared to traditional AT, lightweight defenses like Random Noise Defense (RND) [64] and Attack Against Adversarial [108] (AAA) offer several distinct advantages. They are computationally efficient as they deploy defense during inference, eliminating the need for adversarial training across the entire dataset. This makes the defense easier to deploy on existing models without retraining, offering practical convenience. Furthermore, the flexibility of inference-time-based approaches also allow them to be combined with other defense strategies, such as adversarial training, harnessing the strengths of both. RND and AAA are implemented in BlackboxBench and mainly against query-based attacks. More detailed descriptions of AT, RND, AAA can be found in Appendix E.
III-C Codebase
@@
We build a unified, extensible, modular-based codebase as the basis of BlackboxBench. As shown in Fig. 1, it consists of five modules, including the preparation module (preparing clean data, pretrained model and auxiliary information), attack module, defense module, evaluation module and analysis module.
@@
III-C1 Attack module
The attack module takes preprocessed benign examples, pretrained models and auxiliary information from the preparation module to generate adversarial images. Illustrated in Fig. 4, BlackboxBench summarizes the generation process of various algorithms in decision-based, score-based and transfer-based attacks into three common pipelines, unifying existing methods and illuminate the development of novel attacks. See more details of the unified algorithms in Appendix C. As the pipeline consists of several functional blocks, this attack module is highly extensible to new methods by just modifying individual block functions, without redefining the whole optimization problem. Information regarding how each attack is adapted to this common pipeline is provided in our handbook.
Query-based attack module In the domain of query-based attacks, both score-based and decision-based attack flows can be made up of following four blocks. Nevertheless, they differ significantly inside each block, leading to two independent sub-modules.
@@ III-C1 Attack module
  • Perturbation-generation block In the decision-based attack sub-module, this block computes a perturbation $\textbf{g}_{t+1}$ based on hard-label feedback to narrow the distance between $\boldsymbol{x}_{t}^{*}$ and $\boldsymbol{x}$. In the score-based attack sub-module, it finds a perturbation $\textbf{g}_{t+1}$ based on the probability feedback, aiming to cross the decision boundary within an $\ell_{p}$ norm ball. The employed perturbation generation function is user-optional.
@@
III-C2 Defense module
While our primary focus when building the BlackboxBench lies in implementing black-box attack algorithms, we also offer defense options for users to employ attacks with defense strategies. By integrating this module into the evaluation process, we aim to create a holistic understanding of both the robustness of defensive measures and the potency of attack strategies. This module is designed to rigorously assess the effectiveness of black-box attack algorithms when targeting neural network models equipped with defensive methods. As we mentioned above, we mainly provide 2 types of defense strategies including static defense and dynamic defense in BlackboxBench. Note that the involvement of Adversarial Training defenses is accomplished through the utilization of adversarially trained target models, which is static defense. The defense module exclusively provides three dynamic defense methods: RND[64], AAA-Linear[108], and AAA-Sine[108]. These dynamic defense methods are designed to be lightweight algorithms. Our defense module offers a standardized testing environment, ensuring that evaluations are both consistent and replicable.
III-C3 Evaluation module
The evaluation module includes three crucial metrics: average query number (AQN), median query number (MQN), and attack success rate (ASR). AQN signifies the average query number over all images made during successful query-based black-box attacks. Similarly, MQN captures the middle value among all query numbers of successfully crafted adversarial examples. ASR measures the percent of adversarial samples classified as a different class in untargeted attacks or the target class in targeted attacks. For AQN and MQN, a lower value indicates a better attack performance, while for ASR, a higher value indicates a better attack performance. By employing these three metrics, our module provides a comprehensive evaluation to assess and comprehend black-box adversarial attacks. To best align our evaluation with the process of victim models’ being attacked, we apply an official pre-processing pipeline on adversarial examples to obtain inputs of specific sizes as models may require different sizes of inputs. We note that most prior works didn’t consider transformations during the evaluation.
III-C4 Analysis module
We provide ten analytical tools from four analytical aspects to facilitate our comprehension of black-box adversarial attacks, the details of each tool are presented in Appendix F.
Attack-component effect visualization tools An adversarial attack scenario comprises four fundamental components: the method, attack setting, classification model, and dataset. Nevertheless, the influence of each of these components remains unclear. To unravel the effect of each component, we fix other components, expand the one’s scope, and summarize the results via five plot tools (line charts, scatter charts, bar charts, bubble charts and heatmaps) to explore the underlying patterns.
Adversarial divergence measurement tools BlackboxBench aims to qualitatively interpret the adversarial transferability and establish a quantitative metric for evaluating transferability between white-box surrogate models and the target model. A high transferability implies that the perturbation crafted by the surrogate resembles that crafted by the target. To gain this resemblance, we adopt the Conditional Adversarial Distribution (CAD) modeling technique[18], which models the distribution of perturbations conditioned on clean examples. Based on the CAD, we design a transferability evaluation metric—Adversarial Divergence—to quantify the difference between models from the adversarial standpoint by calculating the normalized Kullback-Leibler divergence between two CADs.
Decision boundary visualization tools To visualize the proximity of benign images to the decision boundary, we employ Decision Surface[120]. We also utilize dbViz[121] to visualize the information of decision boundary around benign images, aiding in understanding the difficulty for attackers trying to misclassify a benign image into other categories than its ground truth.
IV Evaluation and Analysis
Based on the benchmark shown in Sec. III, we further provide thorough evaluations and detailed analyses of implemented black-box adversarial attacks. We first analyze the overview results and attack procedure, then study the effects of data, model architecture, attack budget and defense respectively. Our analyses are broken into several related questions and summarized in Tab. II.
@@ IV Evaluation and Analysis
TABLE II: A summary of analyses in Sec. IV. ● denotes the analyzed question is properly examined in this attack, while ○ indicates it’s not applicable to this attack. Used tool indicates the employed tool from analysis module to facilitate this analysis.
@@
Result overview
1) How has this research community developed?
@@
2) Which attack category exhibits superiority over the others?
AEV
3) Which task is more challenging, targeted or untargeted?
@@
Attack procedure analysis
1) How does the attacker traverse from the benign example to its perturbed counterpart?
@@
Effect of data
1) Whether there are samples harder to be attacked by all attack methods?
@@
2) How does the dimensionality the adversary restricted to affect its performance?
@@
Effect of model architecture
1) Which model architecture is most resistant to malicious queries?
@@
2) How about the transferability between different model architectures?
@@
Effect of attack budget
1) How about the methods’ reliance on the perturbation norm budget?
@@
2) How about the methods’ reliance on the query number budget?
@@
Effect of defense
1) How different attacks perform against various defenses?
@@
2) How different is training-time-based defense versus inference-time-based defense?
AEV
3) Are adversarial images transferable between normally trained models and adversarially trained models?
@@
    AEV $\to$ attack-component effect visualization; MAV-A $\to$ model attention visualization - FullGrad; MAV-B $\to$ model attention visualization - FSM; DBV-A $\to$ decision boundary visualization - Decision Surface; DBV-B $\to$ decision boundary visualization - dbViz; ADM $\to$ adversarial divergence measurement.
TABLE III: Clean accuracies of the models adopted in BlackboxBench. Models in blue are treated as surrogate models. AT indicates adversarially trained models. ${}^{\dagger}$ indicates PyramidNet[53] + ShakeDrop[61] + AutoAugment[62]. VGG-19 ${}^{\dagger}$ indicates VGG-19 with batch normalization.
Query CIFAR-10 ResNet-50[43] ${}^{\dagger}$[44] Inception-V3[45] DenseNet-121[46]
WRN-28-10 (AT w/ DDPM) [65]
WRN-28-10 (AT w/ optimization trick) [2]
93.65% 93.95% 93.74% 94.06%
60.75%
62.80%
ImageNet ResNet-50[43] ${}^{\dagger}$[44] Inception-V3[45] ConVeXt_S[57]
ViT-B/16[55]
ResNet-50 (AT-$\ell_{2}$) [1]
ResNet-50 ($\ell_{\infty}$)[1]
ResNet-50 (FastAT)[91]
76.13% 74.22% 77.29% 83.62%
81.07%
62.83%
54.53%
43.46%
Transfer CIFAR-10 ResNet-50[43] ${}^{\dagger}$[44] Inception-V3[45] DenseNet-BC[46]
ResNeXt-29[49]
WRN-28-10[50]
${}^{\dagger}$[30]
GDAS[54] WRN-28-10 (AT)[60]
95.30% 93.34% 94.77% 96.68%
96.31%
96.21%
98.44%
97.19% 89.69%
ImageNet ResNet-50[43] ${}^{\dagger}$[44] Inception-V3[45] DenseNet-121[46] ViT-B/16[55]
ResNet-152[43]
MobileNet-V2[47]
SENet-154[48] ResNeXt-101[49]
76.15% 74.24% 77.29% 74.65%
81.07%
78.31%
71.88%
81.32% 79.31%
WRN-101[50] PNASNet[51] MNASNet[52] Swin-B[56]
ConvNeXt-B[57]
ResNet-50 (AT)[58]
Swin-B (AT)[59]
ConvNeXt-B (AT)[59]
78.84% 82.90% 73.46% 85.20%
83.80%
64.02%
76.16%
76.02%

### v2

@@ III-A2 Score-based query attacks
The term $\mathbb{B}_{\boldsymbol{x},\epsilon}$ denotes a $\ell_{p}$ norm ball centered at $\boldsymbol{x}$ with radius $\epsilon$. $\xi$ denotes the distance function which returns 0 if $x^{*}$ is located in the ball $\mathbb{B}_{\boldsymbol{x},\epsilon}$ otherwise returns $\infty$. It ensures the imperceptibility of adversarial perturbations. The variable $\Delta$ differentiates based on the type of attack:
@@
Note that the margin loss, captured by the value of $\Delta$, should always be non-negative. A value of 0 for this loss means the attack successfully found its adversarial example, suggesting that the attack process terminates. The primary objective of the attack is to find successful adversarial perturbations while maintaining them within the $\ell_{p}$ norm ball $\mathbb{B}_{\boldsymbol{x},\varepsilon}$.
@@ III-A3 Transfer-based attacks
Unlike query-based attacks that demand repeated queries, transfer-based attacks are inspired by the phenomenon that adversarial samples produced to mislead a surrogate model $f^{\prime}$ can mislead target models $f$, even if their architectures are significantly different [78], as long as $f$ and $f^{\prime}$ solve the same tasks. Let $\mathcal{L}$ be the loss function (e.g. cross-entropy). To fool the classifier $f$, transfer-based attacks use gradient-based methods to craft $\ell_{p}$-norm bounded adversarial perturbation $\boldsymbol{\delta}_{\boldsymbol{x}}$ against the white-box models $f^{\prime}$,
@@
In BlackboxBench, as depicted in Fig. 3, we have implemented a broad range of algorithms for black-box adversarial attacks, spanning the above three types: decision-based query attacks, score-based query attacks, and transfer-based attacks. Each of these types further encompasses different technical routes. We have also implemented various defense methods, including training-time and inference-time approaches. We strongly recommend that readers refer to the Supplementary A for an elaborate taxonomy of the specific algorithms implemented for each category.
III-B Codebase
@@
We build a unified, extensible, modular-based codebase as the basis of BlackboxBench. As shown in Fig. 1, it consists of five modules: preparation module (preparing clean data, pretrained model and auxiliary information), attack module, defense module, evaluation module, and analysis module. Attack module, the core component of the entire codebase, is responsible for generating adversarial samples. Defense module offers options for users to apply defense strategies against attacks. Evaluation module provides metrics to assess the performance of attack methods. Lastly, analysis module incorporates various analytical tools to enhance our understanding of black-box adversarial attacks. Here we only detail the attack module. Supplementary B includes more information about the analysis module, defense module and evaluation module.
@@
III-B1 Attack module
The attack module takes preprocessed benign examples, pretrained models, and auxiliary information from the preparation module to generate adversarial images. As schematically illustrated in Fig. 4, we summarize the generation process of various algorithms in decision-based, score-based and transfer-based attacks into three common pipelines. More detailed unified algorithms can be found in Supplementary F. The attack module of BlackboxBench is constructed based on the unified pipeline. Fig. 34 in the Supplemental Material describes the attack module from an implementation perspective by showing the corresponding component diagrams in Unified Modeling Language (UML).
Query-based attack module For query-based attacks, both score-based and decision-based attack flows can be made up of the following four blocks. Nevertheless, they differ significantly inside each block, leading to two independent submodules.
@@ III-B1 Attack module
  • Perturbation-generation block In the decision-based attack submodule, this block computes a perturbation $\textbf{g}_{t+1}$ based on hard-label feedback to narrow the distance between $\boldsymbol{x}_{t}^{*}$ and $\boldsymbol{x}$. In the score-based attack sub-module, it finds $\textbf{g}_{t+1}$ based on the probability feedback, aiming to cross the decision boundary within an $\ell_{p}$ norm ball. The employed perturbation generation function is user-optional.
@@
By building a unified attack module, BlackboxBench offers several advantages:  1) From the perspective of a fair and standard evaluation, a unified attack module allows differences between methods to exist only in the core function’s code, while the rest of the process remains consistent across all methods. This allows evaluating the method’s actual attack performance.  2) From a researcher’s perspective, BlackboxBench offers inspiration and promising research directions by organizing methods into improvements from different functional blocks within a unified attack process. Meanwhile, this flat procedural programming structure avoids using a large number of inheritance classes, unlike other benchmarks. In this way, code readability is prioritized, ensuring that researchers can quickly understand and begin their work.  3) For engineers, should they wish to implement a more powerful attack by combining multiple attacks, BlackboxBench simplifies integration of each method’s core functional module through defining configuration files. This eliminates need for engineers to redefine the entire optimization process.  4) For contributors, adding a new attack method to BlackboxBench is easy. They simply need to design the core function and register it as the appropriate functional block to integrate it into the unified attack process.
@@
TABLE I: A summary of analyses in Sec. IV. ● denotes the analyzed question is properly examined in this attack, while ○ indicates it’s not applicable to this attack. Used tool indicates the employed tool from analysis module to facilitate this analysis.
@@
2) Which category presents the most promising avenue? AEV
3) Under each attack setting, what is the best-performing method in each category and what is the SOTA method? AEV
@@
2) How different is training-time defense versus inference-time defense? AEV
[…]
@@
    AEV $\to$ attack-component effect visualization (5 tools); MAV-A $\to$ model attention visualization - FullGrad; MAV-B $\to$ model attention visualization - FSM; DBV-A $\to$ decision boundary visualization - Decision Surface; DBV-B $\to$ decision boundary visualization - dbViz; ADM $\to$ adversarial divergence measurement.
IV Evaluation and Analysis
Based on the benchmark shown in Sec. III, we further provide thorough evaluations and detailed analyses of implemented black-box adversarial attacks. In this section, we first analyze the overview results, then study the effects of data, model architecture, and attack budget. We also analyze the influence caused by deployed defense and the attack procedure, please refer to Supplementary C-C and C-D. Our analyses are broken into several related questions, summarized in Tab. I.